OpenStudy (anonymous):

Simplify {the square root of} 2,475 9 sqrt30 11sqrt15 15sqrt11 45sqrt11

OpenStudy (anonymous):

\[\sqrt{2,475}\]

OpenStudy (anonymous):

Help Please !

OpenStudy (anonymous):

good lord

OpenStudy (anonymous):

lol what?

OpenStudy (anonymous):

are you expected to factor that thing??

OpenStudy (anonymous):

lol yes..im expected to.

OpenStudy (anonymous):

if you factor it, you will get \[2475=3^2\times 5^2\times 11\] and so \[\sqrt{2475}=3\times 5\times \sqrt{11}=15\sqrt{11}\] do you know how to factor it?

OpenStudy (anonymous):

Factor..oh im suppose to simplify it lol...i put simplify didn't I?